OpenStudy (anonymous):

y=sin^2(pit) find y' can someone help me please

OpenStudy (anonymous):

I was thinking to use the chain rule so i rewrote the problem ==> (sin(pit)^2 than [2sin(pit)(cospit)

hartnn (hartnn):

yes, and using chain rule only, wouldn't you differentiate pi*x too ? as pi*x is also a function of x.

OpenStudy (kc_kennylau):

f=pi*t, g=sin(f), y=g^2

hartnn (hartnn):

\((\sin^2 (\pi t))' = 2\sin(\pi t)\cos(\pi t)(\pi t)'\)

OpenStudy (anonymous):

how do you differentiate two variables such as (pi*t) ?

OpenStudy (kc_kennylau):

pi is not a variable, it's a constant.

OpenStudy (kc_kennylau):

\(\pi\approx3.14159\)

hartnn (hartnn):

pi, a constant, can be thrown out of differentiation

OpenStudy (anonymous):

so it would go to 0 and i will only have t left?

hartnn (hartnn):

\((\pi t)' = \pi (t)' = \pi\)

OpenStudy (kc_kennylau):

no, you're differentiating with respect to t, so you'll have pi left

hartnn (hartnn):

as derivative of t w.r.t t is 1

OpenStudy (anonymous):

I dont understand how (pit)'=pi

hartnn (hartnn):

what is derivative of x, when differentiated with respect to x ?

OpenStudy (kc_kennylau):

@Anna818 how would you differentiate (3x) wrt x?

OpenStudy (anonymous):

3

hartnn (hartnn):

yes, since pi is also a constant like 3, derivative of pi t will be just pi

OpenStudy (anonymous):

ohh lol okay i got it thanks guys!

OpenStudy (kc_kennylau):

no problem

OpenStudy (anonymous):

From Mathematica 9 Home Edition: \[\frac{\partial \sin ^2(\pi t)}{\partial t}=2 \pi \sin (\pi t) \cos (\pi t) \]

OpenStudy (anonymous):

how would i find the second derivative? i keep getting the wrong answer

hartnn (hartnn):

First derivative = 2pi sin pit cos pit = pi* sin(2pi t) now try

OpenStudy (anonymous):

2cos(pit)(-sinpit)

hartnn (hartnn):

the angle is 2pi*t , isn't it ? and where did the pi(which was outside) go ?

OpenStudy (anonymous):

(2cospit)(pi) (- sinpit) (pi)?

hartnn (hartnn):

\((\pi \sin 2\pi t )' = \pi (\cos(2\pi t)) (2\pi t)' = \pi (\cos(2\pi t)) (2\pi ) =...\)

OpenStudy (anonymous):

where did the second pi come from next to sin ?

hartnn (hartnn):

derivative of 2pi*t 2 pi is constant, so it can be thrown out, = (2 pi) *derivative of t = 2pi *1 =2pi

OpenStudy (anonymous):

so the beginning was 2in(pit) which can be written as (2pit)*2sin?

hartnn (hartnn):

what is "in" in 2in(pit) sin ?

OpenStudy (anonymous):

I dont understand why ur differentiating (pi2sinpit)' 1st differentiating went like this (2sinpit)(cospit)(pi)

OpenStudy (yttrium):

\[y = \sin^2 (\pi t)\]

hartnn (hartnn):

(2sinpit)(cospit)(pi) = pi * (sin (2pi t)) because 2 sin A cos A = sin 2A

OpenStudy (yttrium):

derive using power rule \[y' = n u^{n-1}du\] where u = sin^2 pi t

OpenStudy (yttrium):

hence \[n = 2\] \[u^{n-1} = \sin \pi t\] du = derivative of sin (pit) which is \[du = \pi \cos \pi t\] Hence (by substituting the values from the variables, we'll get \[y' = 2 \sin \pi t (\pi \cos \pi t)\]

OpenStudy (yttrium):

from identity sin2x = 2sinxcosx we can also say thay \[y' = \pi \sin(2 \pi t)\]

OpenStudy (yttrium):

hope this helps.

OpenStudy (anonymous):

I forgot about that identity

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Thanks everyone!

OpenStudy (anonymous):

Yttrium agrees with Mathematica 9 Home Edition.

OpenStudy (anonymous):

here the answer http://www.wolframalpha.com/input/?i=sin%5E2%28pit%29

OpenStudy (anonymous):

2pisin(pit)cos(pit)